OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4

hero (hero):

x =2 -6y - z

OpenStudy (anonymous):

11y + 4z = 6 2y + 5z = 3 y – 2z = 6 –8y – z = 3 3y + 3z = –3 11y + 4z = –6 –9y + z = –3 –y + 2z = –6 THESE ARE THE OPTIONS

hero (hero):

I hate these because they force you to use methods you may not like to use

hero (hero):

-9y + z = -3 -y +2z = -6

OpenStudy (anonymous):

thanks:)